# Iterative Refinement of Project-Level Code Context for Precise Code Generation with Compiler Feedback

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable progress in automated code generation. Yet, incorporating LLM-based code generation into real-life software projects poses challenges, as the generated code may contain errors in API usage, class, data structure, or missing project-specific information. As much of this project-specific context cannot fit into the prompts of LLMs, we must find ways to allow the model to explore the project-level code context. To this end, this paper puts forward a novel approach, termed ProCoder, which iteratively refines the project-level code context for precise code generation, guided by the compiler feedback. In particular, ProCoder first leverages compiler techniques to identify a mismatch between the generated code and the project’s context. It then iteratively aligns and fixes the identified errors using information extracted from the code repository. We integrate ProCoder with two representative LLMs, i.e., GPT-3.5-Turbo and Code Llama (13B), and apply it to Python code generation. Experimental results show that ProCoder significantly improves the vanilla LLMs by over 80% in generating code dependent on project context, and consistently outperforms the existing retrieval-based code generation baselines.
@@ 1 Introduction
Large language models (LLMs), especially those pre-trained on code, as demonstrated by tools such as GitHub Copilot Friedman (2021), Amazon’s CodeWhisperer Amazon (2023), and ChatGPT OpenAI (2023a), are revolutionizing how developers approach programming by automatically generating code based on natural language intent and the context of surrounding code. While existing LLM-based code generation tools excel in handling small code samples, such as completing the current line or the body of a method or giving examples of a standard API, integrating LLM-based code generation into real-life software projects remains challenging Li et al. (2024).
Practical code generation for software projects requires using project-specific contexts like classes, methods, and data structures due to modularity, maintenance, and code independence requirements. However, the repository’s size often exceeds the capacity of the prompt, challenging LLMs to comprehend extensive contexts Liao et al. (2023); Li et al. (2024). Also, studies have demonstrated that irrelevant context in prompts potentially compromises the model performance Shi et al. (2023). Inputting only the task requirements can lead LLMs to overlook project-specific APIs, classes, data structures, or type information specific to a software project’s repository, risking omitting essential logic during code generation Gifany et al. (2013).
Figure 1: Example illustrating limitations of LLM-based code generation: (a) task prompt, (b) wrong solution and error identification, and (c) correct solution utilizing project context.
As a motivation example, consider the case shown in Figure 1. In this case, we use the OpenAI GPT-3.5-Turbo API to generate an image-enhancing function by first reducing the notice and then adjusting brightness and contrast. Given the prompt in Figure 1a, the LLM will produce a code snippet of calling a reduce_noise method as depicted in Figure 1b. Although the generated code follows a standard workflow specified in the task prompt, it leads to a compilation error in our application context because reduce_noise is not implemented. This is unsurprising as the prompt does not provide enough project-level context for the LLM, and can be fixed by providing the project-specific function denoising.denoise as context, as illustrated in Figure 1c. While a carefully engineered prompt may resolve this issue, it is not always possible for the user to generate such prompts, and the project context may be too large to fit into the prompt. As we will show later in the paper, such errors are commonly found when applying LLMs to repository-level code generation Li et al. (2024).
@@
We present ProCoder, a method to allow an LLM to leverage the code repository of a software project to enhance the quality of the generated code. For a given LLM-generated code sample, ProCoder first compiles it and identifies context-related errors. It then retrieves the related context from the code repository to fix the errors. This iterative generation and verification process proceeds repetitively until no error is identified in the generated solution. We demonstrate that ProCoder enhances the accuracy of generation (as indicated by pass rates) by bridging the gap between the repository context and the intended solution.
We evaluate ProCoder by applying it to the CoderEval benchmarking dataset Yu et al. (2023), which consists of code generation tasks utilizing project-specific context. We test ProCoder on two popular code generation models: the GPT-3.5-Turbo OpenAI (2023a) and Code Llama Rozière et al. (2023). Experimental results demonstrate that ProCoder significantly improves the repository-level code generation performance of different dependency levels, outperforming the baseline by over 80% relative pass rates in generating functions dependent on project-specific contexts. Moreover, our iterative method consistently enhances the performance of vanilla retrieval-augmented generation. We also provide a comprehensive analysis of the effectiveness and limitations of ProCoder, offering insights for future research.
@@
  • Extensive experiments and analysis based on two LLMs, i.e., GPT-3.5-Turbo and Code Llama (13B), validate the effectiveness of our proposed ProCoder (Section 5).
@@ 2.1 LLM-based Code Generation
Our work targets the code generation task, which produces source code from a natural-language description complemented by programming context (e.g., project-specific APIs, data structures, and so on). We denote such input as $x$. Given $x$, it is first converted to a sequence of tokens $\bm{x}=[x_{1},\ldots,x_{|\bm{x}|}]$, and a generation model $p_{\text{LM}}(\bm{x})$ predict new tokens sequentially. At each step $t$, the LM calculates the probability distribution of the next token as $p_{\text{LM}}(x_{t}|x_{1:t-1})$. The probability of generating a program $y$ with token sequence $\bm{y}=[x_{|\bm{x}|+1},\ldots,x_{|\bm{x}|+|\bm{y}|}]$ is computed as a product of next-token distributions given left context:
$P(y|x)=\prod_{t=|\bm{x}|+1}^{|\bm{x}|+|\bm{y}|}p_{\text{LM}}(x_{t}|x_{1:t})\,.$ (1)
@@
$P_{\text{LM}}(y|x)=P(y|x,\{\langle x_{i},y_{i}\rangle\}_{i\leq m})\,.$ (2)
Practically, the probability of the next token $x_{t}$ depends on a fixed number of preceding tokens $x_{\max(1,t-w)}:x_{t-1}$, defined by the model’s window length $w$, which may not encompass the entire software project’s code base.
@@ 2.2 Error Analysis in Code Generation
Table 1: Several typical errors reported in compilation and execution.
@@
UNDEF No name ’AsyncBolt5x0’ in module ’neo4j._sync.io._bolt5’
API No value for argument ’xmls’ in function call
OBJECT ’function’ object is not subscriptable
@@
OTHER Parsing failed: ’expected an indented block after function definition
@@
The performance of simple function-level code generation has significantly improved, evidenced by an increase in the pass rate from a pass rate of 31.6% with CodeT5 Wang et al. (2021) to 94.7% in the state-of-the-art Code Llama model Rozière et al. (2023) on the widely-used HumanEval benchmark Chen et al. (2021). However, a recent study Yu et al. (2023) shows that existing LLMs for code generation struggle to generate code snippets that are dependent on the project context information, such as private APIs, types, and variables. To this end, various benchmarks, including ClassEval Du et al. (2023) and CoderEval Yu et al. (2023), have been devised to assess the performance of Language Models in generating context-dependent code within the project.
Figure 2: Distribution of error types in the generated solutions on CoderEval dataset.
Figure 3: Overview of the ProCoder method: (a) the project-level code context extraction process, and (b) iterative refinement to fix compiler-reported errors.
@@
To better illustrate our motivation, we perform an empirical analysis of when the LLMs fail to generate complex code that is dependent on project-level context, on the CoderEval dataset Yu et al. (2023). This dataset comprises 85 function-level tasks and 145 project-level tasks. Specifically, we select the gpt-3.5-turbo OpenAI (2023a) as a target LLM for function-level code generation. Based on it, we choose a state-of-the-art method, termed RepoCoder Zhang et al. (2023a), which retrieves the project-level context as augmentation and feeds 5 code fragments that have the highest DPR similarity score Karpukhin et al. (2020) in prompt for better code generation.
@@
As the code is generated, we compile the code and collect the errors reported by the compiler or in testing. We have generated 10 candidate solutions for a task, comprising 850 solutions for function-level tasks and 1450 for project-level tasks. We report the Pass@10 rate as 53.57% for function-level tasks and 39.73% for project-level tasks. The code that did not pass the test has been taken into error analysis. Each generated code snippet contains precisely one type of error. If multiple errors are reported by the compiler, the most common error type is selected for analysis. The error distribution revealed that four specific types of errors constitute the majority of all errors encountered. We categorize these errors into: 1) UNDEF, involving Use of Undefined Symbol, 2) API, involving Incorrect Use of APIs, 3) OBJECT, involving Improper Use of an Object, 4) FUNC, involving Runtime or Functional Errors, and 5) OTHER, involving Other Syntax and Semantic Errors. Table 1 presents several errors encountered in compiling and testing. One example is the “UNDEF” error, where a variable AsyncBolt5x0 is referenced but does not exist in the specified module.
@@
Figure 2 illustrates the distribution of error types, under the scenario of function-level code generation and project-level code generation. From this figure, we can observe that the majority of errors are runtime or functional errors, accounting for $72\%$ and $34\%$ for function-level and project-level code generation, respectively. Furthermore, the UNDEF errors and the API errors account for substantially high portions of $21\%$ and $64\%$ for function-level and project-level code generation, respectively. ProCoder addresses both types of errors associated with project context by supplying the relevant project context. Experiments demonstrate that ProCoder not only fixes these two types of error but also mitigates other compilation errors by providing feedback on error messages to the language model, leading directly to an improvement in prediction accuracy.
@@ 3.1 Overview
Figure 3 depicts the workflow of ProCoder, consisting of two crucial components: 1) a method for extracting project-level code context through both syntactic and semantic approaches, and 2) a component responsible for iterative generation and evaluation of solutions. This process refines the solutions incrementally, ensuring they evolve towards an error-free state that seamlessly aligns with the project’s environment.
@@ 3.2 Project-Level Code Context Extraction
Supposing that the code generation tools are activated at a specific juncture. In light of the natural language requirement and the code produced by LLMs after an initial iteration, our objective is to extract the semantic context of the generated code from the project’s code base.
Unlike plain texts, source code has syntactic structures that enable precise identification of elements in a project. Thus, in practice, we employ syntax-directed program analysis Harrold and Rothermel (1996) at various points throughout the offline stage to extract the code context at the project level. We initially employ a parser to transform each source code file within the project into an Abstract Syntax Tree (AST), extracting tree nodes that correspond to classes, functions, or variables. Subsequently, if a node of these types is found to be a child of another node (e.g., the function get_handler and the class AsyncBolt in Figure 3(a)’s AST) , an edge is created from the parent node to the child node, establishing a hierarchical relationship.
Take the function $f_{1}$ from Figure 3(a) as an example. From this figure, we can see that both its semantics (e.g., its docstring), and its syntactic relation between the parent class AsyncBolt and file async/bolt.py are captured. This establishes finding the function from the project syntactically using its qualified name async/bolt.py:AsyncBolt.get_handler, or semantically according to its docstring “Return Bolt protocol handlers.”.
@@ 3.3 Retrieval-Augmented Code Generation
Figure 4: Prompt examples for SQL synthesize and code generation.
Based on the compiler feedback, we aim to retrieve the relevant project-level context from all extracted ones. We implement this by transforming the textual compiler feedback into an SQL query using the ChatGPT. The prompt used is presented in Figure 4(a). Here, several examples of paired compiler feedback and SQL queries are provided as demonstrations for in-context learning [^1] Only one demonstration example is illustrated in Figure 3(b), the full list of examples can be found in Appendix C.2..
@@
Using this SQL query, the code snippets that involve the implementations of AsyncBolt3 will be returned from our constructed database. The detailed process of constructing and querying such SQL database is refer to Appendix C.
@@
In addition to returning the project-level level retrieved by the SQL query, we also enhance the acquired context by retrieving similar code snippets using dense passage retrieval. In the initial search round, no compilation error is reported, and ProCoder utilizes the task description string for retrieval. In subsequent searches, ProCoder employs the error report and the corresponding error line.
Given a natural language query $q$, ProCoder first convert it into an embedding vector by utilizing an encoder network, as follows:
$\mathbf{h}_{q}=\text{ENCODER}(q)\,.$ (3)
In our experiments, the pre-trained Transformer Vaswani et al. (2017) is adopted as the encoder network.
After generating the query vector, ProCoder calculate the cosine similarity between the query and embedding vectors of each context entry $\bm{h}_{c}$. This similarity measure is defined as:
$sim(\mathbf{h}_{q},\mathbf{h}_{c})=\dfrac{\mathbf{h}_{q}^{\intercal}\mathbf{h}_{c}}{||\mathbf{h}_{q}||\cdot||\mathbf{h}_{c}||}\,.$ (4)
The top-$n$ entries exhibiting the highest similarity to the query are retrieved as results.
3.4 Refinement with Compiler Feedback
Figure 3(b) showcases the iterative refinement pipeline. Given the task requirement and partial function $f_{1}$, a semantic retrieval is activated to identify similar functions. Specifically, the function $f_{2}$, which provides equivalent functionality in synchronous scenario, is identified. Utilizing both the retrieved context and the prompt illustrated in Figure 4(b), the language model generates a solution; However, the generated output mistakenly invokes SyncBolt3 due to its intended use in asynchronous scenarios, not aligning with the synchronous scenario in $f_{2}$ The compiler’s feedback highlights this error. With this feedback, ProCoder conducts the structural and semantic search, leading to the discovery of the correct function $f_{3}$ for asynchronous scenarios. Incorporating the error details and context into the next iteration ensures accurate function invocation. This process goes iterative until no error is reported by the compiler, resulting in an error-free solution that aligns with the project’s environment. It is noteworthy that ProCoder does not take into account FUNC errors, which arise during execution despite successful compilation. ProCoder focusing on addressing compilation errors, which constitute 66% of the total errors in the context of project-level code generation, as shown in Figure 6.
Table 2: Pass rates of ProCoder based on two LLMs, i.e., GPT-3.5-Turbo and Code Llama (13B), assessed against various baselines across different data splits of CoderEval.
@@ 3.4 Refinement with Compiler Feedback
ProCoder 26.36 39.42 41.82 17.06 29.39 33.82 13.04 28.04 34.78

### v3

@@ Abstract
Large Language Models (LLMs) have shown remarkable progress in automated code generation. Yet, LLM-generated code may contain errors in API usage, class, data structure, or missing project-specific information. As much of this project-specific context cannot fit into the prompts of LLMs, we must find ways to allow the model to explore the project-level code context. We present CoCoGen, a new code generation approach that uses compiler feedback to improve the LLM-generated code. CoCoGen first leverages static analysis to identify mismatches between the generated code and the project’s context. It then iteratively aligns and fixes the identified errors using information extracted from the code repository. We integrate CoCoGen with two representative LLMs, i.e., GPT-3.5-Turbo and Code Llama (13B), and apply it to Python code generation. Experimental results show that CoCoGen significantly improves the vanilla LLMs by over 80% in generating code dependent on the project context and consistently outperforms the existing retrieval-based code generation baselines.
@@ 1 Introduction
Large Language Models (LLMs), especially those pre-trained on code, as demonstrated by tools such as GitHub Copilot Microsoft (2024), Amazon’s CodeWhisperer Amazon (2023), and ChatGPT OpenAI (2023), are revolutionizing how developers approach programming by automatically generating code for given contexts (e.g., natural-language descriptions or surrounding incomplete code). While existing LLM-based code generation tools excel in code generation within a small-scale and isolated context (i.e., a single file), integrating LLM-based code generation into real-world software projects remains challenging Li et al. (2024).
Practical code generation for software repositories is often associated with broader project-level contexts declared in other repository files due to the demands of modularity, structure, and comprehension in software management Kemerer (1995). Solely providing the task requirements and surrounding incomplete code can lead LLMs to overlook the complex hierarchies of APIs, classes, data structures, or type constraints specific to a software project’s repository, risking omitting essential logic during code generation Gifany et al. (2013).
Figure 1: LLM-based code generation example. (a) task prompt; (b) wrong solution and error identification; (c) correct solution utilizing project context
As a motivating example, consider the case shown in Figure 1. In this case, we use the OpenAI GPT-3.5-Turbo API to generate an image-enhancing function by first reducing the notice and then adjusting brightness and contrast. Given the prompt in Figure 1a, the LLM will produce a code snippet of calling a reduce_noise method as depicted in Figure 1b. Although the generated code follows a standard workflow specified in the task prompt, it leads to a compilation error in our application context because reduce_noise is not implemented. This is unsurprising as the prompt does not provide enough project-level context for the LLM, and can be fixed by providing the project-specific function denoising.denoise as context, as illustrated in Figure 1c. While a carefully engineered prompt may resolve this issue, it is not always possible for the user to generate such prompts, and the project context may be too large to fit into the prompt. As we will empirically show later in (Section 1), such errors are commonly found when applying LLMs to repository-level code generation Li et al. (2024).
@@
We present CoCoGen, a method to allow an LLM to leverage the code repository of a software project to enhance the quality of the generated code. For a given LLM-generated code sample, CoCoGen first compiles it and identifies context-related errors. It then retrieves the related context from the code repository to fix the errors. This iterative generation and verification process proceeds repetitively until no error is identified in the generated solution. We demonstrate that CoCoGen enhances the accuracy of generation (as indicated by pass rates) by bridging the gap between the repository context and the intended solution.
We evaluate CoCoGen by applying it to the CoderEval benchmarking dataset Yu et al. (2024), which consists of code generation tasks utilizing project-specific context. We test CoCoGen on two popular code generation models: the GPT-3.5-Turbo OpenAI (2023) and Code Llama Rozière et al. (2023). Experimental results demonstrate that CoCoGen significantly improves the repository-level code generation performance of different dependency levels, outperforming the baseline by over 80% relative pass rates in generating functions dependent on project-specific contexts. Moreover, our iterative method consistently enhances the performance of vanilla retrieval-augmented generation. We also provide a comprehensive analysis of the effectiveness and limitations of CoCoGen, offering insights for future research.
@@
  • Extensive experiments and analysis based on two LLMs, i.e., GPT-3.5-Turbo and Code Llama (13B), showing the effectiveness of the proposed CoCoGen method (Section 5).
The source code and dataset used in this paper are available at: https://github.com/CGCL-codes/naturalcc/tree/main/examples/cocogen.
@@ 2.1 LLM-based Code Generation
Our work targets the code generation task, which produces source code from a natural-language description complemented by programming context (e.g., project-specific APIs, and data structures). We denote this input as $x$. Given $x$, it is first converted to a sequence of tokens $\bm{x}=[x_{1},\ldots,x_{|\bm{x}|}]$, and a generative Language Model (LM) $p_{\text{LM}}(\bm{x})$ predicts new tokens sequentially. At each step $t$, the LM calculates the probability distribution of the next token as $p_{\text{LM}}(x_{t}|x_{1:t-1})$. The probability of generating a program $y$ with token sequence $\bm{y}=[x_{|\bm{x}|+1},\ldots,x_{|\bm{x}|+|\bm{y}|}]$ is computed as a product of next-token distributions given left context:
$p(y|x)=\prod_{t=|\bm{x}|+1}^{|\bm{x}|+|\bm{y}|}p_{\text{LM}}(x_{t}|x_{1:t})$ (1)
@@
$p_{\text{LM}}(y|x)=p(y|x,\{\langle x_{i},y_{i}\rangle\}_{i\leq m})$ (2)
Practically, the probability of the next token $x_{t}$ depends on a fixed number of preceding tokens $x_{\max(1,t-w)}:x_{t-1}$, defined by the model’s context window length $w$, without encompassing the entire software project’s code base.
@@ 2.2 Error Analysis in Code Generation
Table 1: Typical errors reported in compilation and execution
@@
UNDEF No name ‘AsyncBolt5x0’ in module ’neo4j._sync.io._bolt5’
API No value for argument ‘xmls’ in function call
OBJECT ‘function’ object is not subscriptable
@@
OTHER Parsing failed: ‘expected an indented block after function definition’
@@
The performance of simple function-level code generation has significantly improved, as demonstrated by an increase in the pass rate from 31.6% with CodeT5 Wang et al. (2021) to 94.7% with the state-of-the-art Code Llama Rozière et al. (2023) on the widely-used HumanEval benchmark Chen et al. (2021). However, a recent study Yu et al. (2024) shows that existing LLMs for code generation struggle to generate code snippets that are dependent on the project contexts, such as private APIs, classes, data structures, or type constraints. To this end, various benchmarks, including ClassEval Du et al. (2024), CoderEval Yu et al. (2024), and CrossCodeEval Ding et al. (2023), have been devised to assess the performance of LLMs in generating context-dependent code within the project.
@@
To better illustrate our motivation, we perform an empirical analysis of when the LLMs fail to generate complex code that is dependent on project-level context, on the CoderEval dataset Yu et al. (2024). This dataset comprises 85 function-level tasks and 145 project-level tasks. Specifically, we select the GPT-3.5-Turbo OpenAI (2023) as a target LLM for function-level code generation. Based on it, we select a state-of-the-art method called RepoCoder Zhang et al. (2023a). RepoCoder retrieves project-level context as an augmentation and incorporates the five code fragments with the highest similarity scores, as determined by dense passage retrieval Karpukhin et al. (2020), into the prompt for improved code generation.
Figure 2: Distribution of error types in the generated solutions on CoderEval dataset
Figure 3: Overview of the CoCoGen method. (a) the project-level code context extraction process; (b) iterative refinement to fix compiler-reported errors
@@
As the code is generated, we compile it and collect any errors reported by the compiler or encountered during testing. We have generated 10 candidate solutions for a task, comprising 850 solutions for function-level tasks and 1450 for project-level tasks. We report the Pass@10 rate as 53.57% for function-level tasks and 39.73% for project-level tasks. The code that does not pass the test has been taken into error analysis. Each generated code snippet contains precisely one type of error. If multiple errors are reported by the compiler, the most common error type is selected for analysis. The error distribution reveals that four specific types of errors constitute the majority of all errors encountered. We categorize these errors into: 1) UNDEF, involving Use of Undefined Symbol, 2) API, involving Incorrect Use of APIs, 3) OBJECT, involving Improper Use of an Object, 4) FUNC, involving Runtime or Functional Errors, and 5) OTHER, involving Other Syntax and Semantic Errors. Table 1 presents several errors encountered in compiling and testing. One example is the “UNDEF” error, where a variable AsyncBolt5x0 is referenced but does not exist in the specified module.
@@
Figure 2 illustrates the distribution of error types, under the scenario of function-level code generation and project-level code generation. From this figure, we can observe that the majority of errors are runtime or functional errors, accounting for $72\%$ and $34\%$ for function-level and project-level code generation, respectively. Furthermore, the UNDEF errors and the API errors account for substantially high portions of $21\%$ and $64\%$ for function-level and project-level code generation, respectively. CoCoGen addresses both types of errors associated with project context by supplying the relevant project context. Experiments demonstrate that CoCoGen not only fixes these two types of error but also mitigates other compilation errors by providing context feedback on error messages to the code LM, leading directly to an improvement in prediction accuracy.
@@ 3.1 Overview
Figure 3 depicts the workflow of CoCoGen, consisting of two crucial components: 1) a method for extracting project-level code context through both syntactic and semantic approaches, and 2) a component responsible for iterative generation and evaluation of solutions. This process refines the generated solutions incrementally, ensuring they evolve towards an error-free state that seamlessly aligns with the codebase of the software project.
@@ 3.2 Project-Level Code Context Extraction
Supposing that the code generation tools are activated at a specific juncture. In light of the natural-language requirement and the code produced by LLMs after an initial iteration, our objective is to extract the semantic context of the generated code from the project’s code base.
Unlike plain texts, source code has syntactic structures that enable precise identification of elements in a project. Thus, in practice, we employ syntax-directed program analysis Aho et al. (2006) at various points throughout the offline stage to extract the code context at the project level. We initially employ a parser to transform each source code file within the project into an Abstract Syntax Tree (AST), extracting tree nodes that correspond to classes, functions, or variables. Subsequently, if a node of these types is found to be a child of another node (e.g., the function get_handler and the class AsyncBolt in Figure 3a’s AST) , an edge is created from the parent node to the child node, establishing a hierarchical relationship.
Take the function $f_{1}$ from Figure 3a as an example. From this figure, we can see that both its semantics (e.g., its docstring), and its syntactic relation between the parent class AsyncBolt and file async/bolt.py are captured. This allows CoCoGen to find the function from the project syntactically using the function’s qualified name AsyncBolt.get_handler(), or semantically according to its docstring “Return Bolt protocol handlers”.
@@ 3.3 Retrieval-Augmented Code Generation
Figure 4: Prompt examples for (a) SQL synthesize and (b) code generation
Structural Search.
Based on the compiler feedback, we aim to retrieve the relevant project-level context from all extracted ones. We implement this by transforming the textual compiler feedback into an SQL query using the ChatGPT. The prompt used is presented in Figure 4a. Several examples of paired compiler feedback and SQL queries are provided as demonstrations for in-context learning***Only one demonstration example is illustrated in Figure 3(b), the full list of examples can be found in Appendix G.2..
@@
Using this SQL query, the code snippets that involve the implementations of AsyncBolt3 will be returned from our constructed database. The detailed process of constructing and querying such SQL database is refer to Appendix G.
@@
In addition to returning the project-level context retrieved by the SQL query, we also enhance the acquired context by retrieving similar code snippets using dense passage retrieval Karpukhin et al. (2020). In the initial search round, no compilation error is reported, and CoCoGen utilizes the task description string for retrieval. In subsequent searches, CoCoGen employs the error report and the corresponding error line.
Given a natural-language query $q$, CoCoGen first converts it into an embedding vector by utilizing an encoder network, as follows:
$\mathbf{h}_{q}=\text{ENCODER}(q)$ (3)
CoCoGen utilizes a pre-trained Transformer network Vaswani et al. (2017) as the encoder. After generating the query vector, CoCoGen calculates the cosine similarity between the query and embedding vectors of each context entry $\bm{h}_{c}$. This similarity measure is defined as:
$sim(\mathbf{h}_{q},\mathbf{h}_{c})=\dfrac{\mathbf{h}_{q}^{\intercal}\mathbf{h}_{c}}{||\mathbf{h}_{q}||\cdot||\mathbf{h}_{c}||}$ (4)
and the top-$n$ entries exhibiting the highest similarity to the query are retrieved as results.
Table 2: Pass rates of CoCoGen based on two LLMs, i.e., GPT-3.5-Turbo and Code Llama (13B), assessed against various baselines across different splits of the CoderEval dataset
@@
CoCoGen 26.36 39.42 41.82 17.06 29.39 33.82 13.04 28.04 34.78
3.4 Refinement with Compiler Feedback
Figure 3(b) showcases the iterative refinement pipeline. Given the task requirement and partial function $f_{1}$, a semantic retrieval is activated to identify similar functions. Specifically, the function $f_{2}$, which provides equivalent functionality in synchronous scenario, is identified. Utilizing both the retrieved context and the prompt illustrated in Figure 4(b), the language model generates a solution.
However, the generated output mistakenly invokes SyncBolt3 due to its intended use in asynchronous scenarios, not aligning with the synchronous scenario in $f_{2}$ The compiler’s feedback highlights this error. With this feedback, CoCoGen conducts the structural and semantic search, leading to the discovery of the correct function $f_{3}$ for asynchronous scenarios. Incorporating the error details and context into the next iteration ensures accurate function invocation. This process goes iterative until no error is reported by the compiler, resulting in an error-free solution that aligns with the project’s environment.
It is noteworthy that CoCoGen does not take into account FUNC errors, which arise during execution despite successful compilation. CoCoGen focuses on addressing compilation errors, which constitute 66% of the total errors in the context of project-level code generation, as shown in Figure 6.